OpenStudy (anonymous):

Identify the 19th term of the arithmetic sequence 5, 16, 27 …

OpenStudy (anonymous):

an=a1 + (n-1)d a19 = 5 +(19-1) ... then I'm stuck

OpenStudy (mertsj):

What is the common difference?

OpenStudy (mertsj):

5+x=16 16+x=27 27+x= the fourth term What is x ?

OpenStudy (anonymous):

11

OpenStudy (mertsj):

So replace d in your formula with 11 and complete the calculation

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

a19=203?

OpenStudy (mertsj):

Yep. That's what I got as well.

OpenStudy (anonymous):

yay thanks (:

OpenStudy (anonymous):

what if its fractions

OpenStudy (mertsj):

The process is the same. Do you have an example?

OpenStudy (anonymous):

Identify the 25th term of the arithmetic sequence 2, 1(3/5), 1(1/5)

OpenStudy (anonymous):

so its going by 2/5?

OpenStudy (mertsj):

Do you mean 1times 3/5 or 1 and 3/5?

OpenStudy (anonymous):

1 and 3/5

OpenStudy (mertsj):

2+x=1 3/5 1 3/5+x= 1 1/5 1 2/5 + x = the fourth term What is x?

OpenStudy (anonymous):

i don't know how to find that

OpenStudy (mertsj):